OpenStudy (anonymous):

verify if the following measures are side of a right triangle 1.) 3,4,5 2.) 5,12,13 3.) 6,8,10 4.) 1,1,2 5.)4,4,5

OpenStudy (mertsj):

Use the Pythagorean Theorem Does 3^2+4^2=5^5 9+16=25 yes! these are the measures of the sides of a right triangle